OpenStudy (anonymous):

What is the limit as x approaches -3 of the function 3-lxl/ 3+x

OpenStudy (anonymous):

Graph it and see what happens.

OpenStudy (anonymous):

im not allowed to use a graph on this one

OpenStudy (anonymous):

What are you allowed to use?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%283-+absolute+value+of+x%29%2F+%283%2Bx%29 here it's reaching 0 but when i do it algebraically i get 1

OpenStudy (anonymous):

I'm allowed to do it algebraically only lol @wio

OpenStudy (anonymous):

Okay, if you want to do it algebraically, then you have to split the function up into two cases: When \(x\) is negative and when it is positive.

OpenStudy (anonymous):

oh i got it i as looking at positive 3 thank you anyway

OpenStudy (anonymous):

\[ f(x) =\begin{cases} \frac{3-x}{3+x} &x>0\\ \frac{3+x}{3+x} &x<0 \end{cases} \]We want to use that \(x<0\) case.

OpenStudy (anonymous):

\[ \lim_{x\to -3}\frac{3-|x|}{3+x} = \lim_{x\to -3}\frac{3+x}{3+x} = \lim_{x\to -3}1 = 1 \]

OpenStudy (anonymous):

No medal?